OpenStudy (anonymous):

how do solve this? 64-9x^2=0

OpenStudy (anonymous):

would it be difference of 2 squares?

OpenStudy (anonymous):

x^2=64/9 x=+-8/3

OpenStudy (anonymous):

i dont think thats it...

OpenStudy (experimentx):

i think thats it ..

OpenStudy (anonymous):

Difference of two squares, yes.